# PennyLang: Pioneering LLM-Based Quantum Code Generation with a Novel PennyLane-Centric Dataset

### v2

@@ Abstract
Large Language Models (LLMs) offer remarkable capabilities in code generation, natural language processing, and domain-specific reasoning. However, their application in quantum software development remains underexplored, particularly for PennyLane—a leading framework for hybrid quantum-classical computing. To address this gap, we introduce a novel, high-quality dataset comprising 3,347 PennyLane-specific quantum code samples and contextual descriptions, specifically curated to support LLM training and fine-tuning for quantum code assistance. Our contributions are threefold: (1) the automatic construction and open-source release of a comprehensive PennyLane dataset derived from textbooks, official documentation, and open-source repositories; (2) a structured methodology for data curation, annotation, and formatting to enhance LLM usability and relevance; and (3) a rigorous evaluation of code generation capabilities using both baseline Retrieval-Augmented Generation (RAG) and a GraphRAG-enhanced pipeline. Using the PennyLang framework, we demonstrate that GraphRAG, when applied to a GPT-4o Mini model, substantially outperforms standard prompting and baseline RAG. Accuracy improves from 20.5% (without RAG) to 58.2% with GraphRAG, showcasing its effectiveness in reducing hallucinations and improving code correctness in quantum programming tasks. Compared to prior efforts focused largely on Qiskit, our work expands LLM-based assistance to the PennyLane ecosystem, contributing practical tools and reproducible methodologies for advancing AI-assisted quantum software development.
@@ I Introduction
Quantum computing has emerged as a transformative technology with the potential to solve complex problems that are infeasible for classical systems [1, 2, 3]. Breakthroughs such as Google’s demonstration of “quantum supremacy” [4]—where a quantum processor performed a calculation in seconds that would take thousands of years on a classical supercomputer—highlight its disruptive capabilities in domains ranging from cryptography and quantum chemistry to machine learning and optimization [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15].
Research Challenges: Despite these advances, programming quantum systems is challenging due to the specialized nature of existing frameworks and benchmarks, hindering widespread adoption [16, 17, 18]. PennyLane is a leading open-source python framework designed for hybrid quantum-classical computing, allowing seamless integration of quantum circuits with machine learning workflows [19]. However, unlike the Qiskit framework [20, 21], which benefits from dedicated AI-driven code assistants [22], PennyLane lacks equivalent tools to assist developers—particularly newcomers—in writing optimized quantum code.
In parallel, the field of Natural Language Processing (NLP) has seen rapid advancements in Large Language Models (LLMs), which now excel at reasoning, contextual understanding, and automated code generation [23, 24, 25]. These models have shown success in classical programming [26, 27], improving developer productivity and reducing syntax and logic errors. However, their application in quantum programming—especially within the PennyLane ecosystem—remains underexplored. This motivates our work: curating a high-quality, PennyLane-centric dataset and establishing a structured methodology for refinement, annotation, and evaluation, ultimately enabling LLM-based quantum code generation and debugging.
State-of-the-Art and Limitations: While LLMs have been benchmarked extensively for classical programming [28, 29], their performance in quantum domains is less understood. Qiskit HumanEval and Qiskit Code Assistant [22] represent early attempts to extend LLM capabilities into quantum programming, but they focus primarily on Qiskit—a hardware-centric framework. PennyLane, in contrast, is designed for variational quantum algorithms and quantum machine learning [30], requiring a different kind of LLM support and training corpus.
The primary challenge in developing a PennyLane code assistant lies in data availability. Unlike classical programming languages, where vast repositories of well-documented code exist, quantum computing resources are fragmented across research papers, user forums, open-source repositories, and official documentation [31]. Moreover, raw quantum code samples often lack contextual annotations, making it difficult for LLMs to learn meaningful associations between quantum operations and their intended applications.
Our Approach and Contributions: In this work, we introduce PennyLang, a large-scale, high-quality dataset tailored for PennyLane code generation, along with a rigorous evaluation pipeline. In addition to a baseline RAG framework, we propose and test the GraphRAG framework [32], which leverages structured code relationships to improve the quality and relevance of retrieval during generation. Our contributions are summarized in Fig. 1 and detailed below:
@@
    Comprehensive PennyLane Dataset: We curate a dataset of 3,347 PennyLane-specific code samples sourced from GitHub repositories, quantum computing textbooks, and official PennyLane documentation. Each sample is enriched with contextual descriptions and annotations to support instruction tuning.
@@
    Systematic Data Refinement and Annotation: We present a robust methodology for dataset cleaning and metadata enrichment. Using the GPT-4o API, code snippets are transformed into instruction-query pairs, improving compatibility with LLM training pipelines.
@@
    RAG-Based Evaluation Framework: We use a Retrieval-Augmented Generation (RAG) evaluation framework to assess the dataset’s impact on model performance. The framework compares outputs from three models—GPT-4o mini, Claude 3.5, and Qwen 7B—under two conditions: (i) with access to the dataset as retrieval-augmented context and (ii) without dataset augmentation. The generated outputs are validated and scored based on functionality, syntax correctness, and modularity.
  4. 4.
    GraphRAG-Enhanced Evaluation Framework: We extend the baseline RAG framework with GraphRAG, which introduces a graph-based context retrieval mechanism for code augmentation. GraphRAG models contextual and structural relationships between quantum code samples, enabling more accurate and consistent retrievals. We evaluate this setup using GPT-4o Mini model, comparing GraphRAG, and no-RAG scenarios in a controlled setting.
Summary of Key Results: RAG-enhanced quantum code generation improves performance across all models. GPT-4o Mini showed an 11.67% gain in accuracy and modularity, Claude 3.5 Sonnet improved by 7.6%, enhancing coherence, while Qwen 2.5 saw the highest boost at 14.38%, benefiting from retrieved context. These results confirm RAG’s effectiveness in optimizing PennyLane quantum code generation by improving functionality, syntax, and adaptability. Our results from GraphRAG significantly outperforms both baseline RAG and vanilla prompting. When applied to GPT-4o Mini, GraphRAG boosts accuracy from 20.5% (no RAG) to 58.2%, marking a 37.7% absolute improvement in correct solutions across PennyLang dataset, confirming the benefit of structured retrieval in quantum programming tasks.
Paper Organization: Sec. II reviews related work and background on quantum programming and RAG. Sec. III outlines our dataset construction, annotation methodology, and the GraphRAG retrieval pipeline. Sec. IV presents experimental results comparing model performance across prompting settings. Sec. V discusses case studies and qualitative insights. Sec. VI concludes with future directions for dataset-driven quantum code assistance.
@@ II Background and Related Work
TABLE I: Summary of Qiskit and PennyLane Datasets for LLM Training.
Dataset Name Framework Source Samples Available Dataset Type Format Qiskit Textbook Dataset [33] Qiskit Qiskit.org Textbook Multiple structured examples Structured Qiskit Code Python Scripts IBM Quantum Challenge Data [34] Qiskit IBM Quantum Experience Derived from past challenges Instruction-Response Format JSON/Notebook Qiskit Code Repository [35] Qiskit GitHub (Qiskit AI repos) Thousands (Scattered) Raw Qiskit Code Samples Python Scripts Q-Gen Quantum Circuit [36] Qiskit Kaggle (Possible) 309 (Limited samples) Pretrained Benchmark Data CSV/JSON QDataSet [37] Qiskit GitHub 52 (raw datasets) Quantum ML Data Various QCircuitNet [38] Qiskit Research Paper 4,800 Samples Quantum Algorithm Design Benchmark Python/QASM PennyLane Documentation [39] PennyLane PennyLane.ai Docs Various examples Raw PennyLane Code Samples Python Scripts PennyLane GitHub Repos [40] PennyLane GitHub Open-Source Scattered code samples Unstructured Quantum Code Python Scripts PennyLang (Ours) PennyLane Open-Source Dataset 3,347 Instruction-Response Dataset Curated and Structured Dataset JSON file
The intersection of LLMs and quantum computing is an emerging research domain. While LLMs have demonstrated impressive capabilities in NLP [41] and automated code generation [42], their application in quantum software development remains underexplored. Most existing studies focus on classical programming languages or quantum frameworks such as Qiskit [22], with little attention to the PennyLane framework [19], despite its popularity in hybrid quantum-classical workflows.
This section presents an in-depth review of related work in three key areas:
@@ II-A LLMs for Code Generation in Classical Programming
Transformer-based models have significantly impacted AI-driven programming, automating code generation, completion, and debugging. Key contributions include:
  • Codex (OpenAI) [26] introduced a language model trained on public code repositories, enabling natural language-to-code translation.
  • CodeBERT [43] leveraged bidirectional transformers to enhance source code understanding.
  • InCoder [29] developed a multi-turn interactive model for code infilling and synthesis.
  • AlphaCode [25] showed competitive programming-level LLM capabilities.
Additionally, benchmarking studies have evaluated LLM performance on programming tasks:
  • HumanEval [24] proposed an execution-based evaluation for Python code generation.
  • MBPP [44] introduced a multi-language benchmark covering correctness and execution constraints.
  • CodeXGLUE [45] aggregated multiple datasets to benchmark code-related tasks.
These studies highlight the necessity of well-structured datasets to enhance LLM code generation performance.
@@ II-B LLMs for Quantum Programming
Existing research on AI-powered quantum code assistants focuses primarily on the Qiskit framework:
  • Qiskit Code Assistant (IBM Research) [22] assists users in debugging and optimizing Qiskit code.
  • Qiuchi et al. [46] adapted BERT-based models to improve quantum code understanding.
  • CodeLlama-Quantum (Meta AI) [27] extended CodeLlama for Qiskit-based applications.
However, these works face key limitations:
  1. 1.
    Lack of High-Quality Data: Quantum programming lacks scalable, curated training datasets.
  2. 2.
    Code Complexity: Quantum circuits require LLMs to reason over non-trivial entanglement and measurement logic.
  3. 3.
    Hybrid Workflows: PennyLane introduces mixed classical-quantum code patterns that require deeper contextual understanding.
To bridge this gap, we propose the first PennyLane-centric dataset—PennyLang—paired with an advanced evaluation pipeline that incorporates RAG-based retrieval.
II-C Retrieval-Augmented Generation and GraphRAG
Retrieval-Augmented Generation (RAG) [47] has emerged as a popular method to ground LLM outputs using external context. In quantum programming, RAG helps models compensate for limited training data and outdated knowledge by retrieving relevant examples at inference time.
However, baseline RAG systems suffer from semantic drift, irrelevant context injection, and weak code relevance due to shallow vector-based similarity matching. To address this, we employ GraphRAG[32], a graph-enhanced retrieval framework designed to capture structural and functional relationships between quantum code samples. GraphRAG models PennyLang as a graph of interconnected examples, allowing retrievals that are topologically and semantically aligned with the user query.
Compared to traditional RAG, GraphRAG significantly improves context precision and reduces hallucinations in generated code. This leads to large gains in generation accuracy on the PennyLang test dataset.
Figure 2: Methodology for PennyLang quantum code generation and evaluation. Data is collected from GitHub, quantum books, and PennyLane docs, then refined, tokenized, and formatted for LLMs. The curated dataset is processed by GPT-4o for input-output conversion. User queries are handled via a retrieval-augmented generation (RAG) framework using LangChain and Chroma to fetch relevant code. Responses from GPT-4o Mini, Claude 3.5 Sonnet, and Qwen2.5-7B-Instruct-Turbo are compared against non-RAG outputs and evaluated on functionality, syntax, and modularity to assess RAG’s impact.
Our contributions advance AI-assisted quantum programming by enabling LLMs to provide high-quality, domain-specific assistance in PennyLane.
TABLE II: Comparison of Related Work and Our Contributions
Research Area Existing Works Our Contributions.
LLMs for Code Generation Codex [26], CodeBERT [43] PennyLane-specific dataset
LLMs for Quantum Programming Qiskit Assistant [22], Qiuchi et al. [46] Hybrid quantum-classical dataset
Quantum Code Datasets QCircuitNet [38], QDataset  [38] First PennyLane dataset
@@ III-A1 Data Collection from GitHub Repositories
To systematically gather PennyLane-specific code, we utilized the GitHub API to search for publicly accessible repositories with permissive open-source licenses (e.g., MIT, Apache 2.0, BSD). The selection criteria required that the term "PennyLane" be present in either the repository name, description, or README file.
To maintain dataset quality and avoid redundancy, we followed these constraints:
@@
  • Only .py files (Python scripts) were extracted for further analysis.
@@
Since PennyLane maintains backward compatibility and rarely deprecates features [19], we did not impose a time constraint on data collection, ensuring a broad dataset scope. Additionally, introductory comment blocks containing license information, author details, or unrelated metadata were removed to retain only executable and relevant code.
This curation process resulted in a high-fidelity dataset of PennyLane-specific implementations. In total, 1,321 samples were collected from the official PennyLaneAI GitHub repository [40].
@@ III-A2 Extracting Quantum Code from Books
In addition to GitHub, we extracted PennyLane-related code snippets from two quantum computing books:
  • “Quantum Machine Learning: An Applied Approach” [48]
  • “A Practical Guide to Quantum Machine Learning and Quantum Optimization” [49]
These books provide theoretical foundations and applied examples of PennyLane-based quantum machine learning (QML) and variational quantum algorithms (VQAs). The extraction process involved:
@@
    Manually verification of each snippet with explanations from the books, converting descriptive text into structured comments.
@@ III-A3 Scraping Code from PennyLane Documentation
The official PennyLane documentation [39] serves as a comprehensive reference for quantum programming, covering:
  • Quantum circuits and variational algorithms.
  • Quantum operators and templates.
  • Hybrid quantum-classical optimization.
  • Quantum chemistry applications.
TABLE III: Tokenized Input and Attention Masking.
Tokenized Input Attention Mask
"<pad>" 0
"<pad>" 0
"def" 1
":" 1
"circuit" 1
Tokenized Input Attention Mask
"(" 1
")" 1
"return" 1
"qml.expval" 1
To further enhance our dataset, we systematically scraped all code snippets from the official PennyLane website. This included examples embedded in tutorials and API references.
As with book sources, documentation explanations were manually converted into comments, ensuring that each code sample contained self-contained, contextualized information. This practice enhances LLM interpretability, as models trained on such data can generate both code and explanations.

### v3

@@ Abstract
Large Language Models (LLMs) offer powerful capabilities in code generation, natural language understanding, and domain-specific reasoning. Their application to quantum software development remains limited, in part because of the lack of high-quality datasets both for LLM training and as dependable knowledge sources. To bridge this gap, we introduce PennyLang, an off-the-shelf, high-quality dataset of 3,347 PennyLane-specific quantum code samples with contextual descriptions, curated from textbooks, official documentation, and open-source repositories. Our contributions are threefold: (1) the creation and open-source release of PennyLang, a purpose-built dataset for quantum programming with PennyLane; (2) a framework for automated quantum code dataset construction that systematizes curation, annotation, and formatting to maximize downstream LLM usability; and (3) a baseline evaluation of the dataset across multiple open-source models, including ablation studies, all conducted within a retrieval-augmented generation (RAG) pipeline. Using PennyLang with RAG substantially improves performance: for example, Qwen 7B’s success rate rises from 8.7% without retrieval to 41.7% with full-context augmentation, and LLaMa 4 improves from 78.8% to 84.8%, while also reducing hallucinations and enhancing quantum code correctness. Moving beyond Qiskit-focused studies, we bring LLM-based tools and reproducible methods to PennyLane for advancing AI-assisted quantum development.
@@ I Introduction
Quantum computing has emerged as a transformative technology with the potential to solve complex problems that are infeasible for classical systems [1, 2, 3]. Breakthroughs such as Google’s demonstration of “quantum supremacy” [4], where a quantum processor performed a calculation in seconds that would take thousands of years on a classical supercomputer, highlight its disruptive capabilities in domains ranging from cryptography and quantum chemistry to machine learning and optimization [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16].
Research Challenges: Despite these advances, programming quantum systems is challenging due to the specialized nature of existing frameworks and benchmarks, hindering widespread adoption [17, 18]. PennyLane is a leading open-source Python framework designed for hybrid quantum-classical computing, allowing seamless integration of quantum circuits with machine learning workflows [19]. However, unlike the Qiskit framework [20, 21], which benefits from dedicated AI-driven code assistants [22], PennyLane lacks equivalent tools to assist developers in writing optimized quantum code.
In parallel, the field of Natural Language Processing (NLP) has seen rapid advancements in Large Language Models (LLMs), which now excel at reasoning, contextual understanding, and automated code generation [23, 24, 25]. These models have shown success in classical programming [26, 27], improving developer productivity and reducing syntax and logic errors. However, their application in quantum programming, especially within the PennyLane ecosystem, remains underexplored. This motivates our work: curating a high-quality, PennyLane-centric dataset and establishing a structured methodology for refinement, annotation, and evaluation, ultimately enabling LLM-based quantum code generation and debugging.
State-of-the-Art and Limitations: LLMs have been extensively benchmarked for classical programming [28, 29], but their performance in quantum domains remains less explored. Early efforts such as Qiskit HumanEval and Qiskit Code Assistant [22] adapt LLMs to quantum programming but focus on Qiskit’s hardware-centric paradigm. PennyLane, by contrast, targets variational quantum algorithms and quantum machine learning [30], requiring dedicated LLM support and a specialized training corpus. A major barrier is the scarcity of suitable data: quantum code is scattered across papers, forums, repositories, and documentation [31], and often lacks the contextual annotations necessary for LLMs to link quantum operations with their intended use.
@@
Our Approach and Contributions: In this work, we introduce PennyLang, a large-scale, high-quality dataset specifically designed for PennyLane code generation, accompanied by a rigorous evaluation pipeline. We implement a baseline Retrieval-Augmented Generation (RAG) framework to assess the dataset’s effectiveness across multiple models. Our key contributions, shown in Fig. 1, consist of:
@@
    Comprehensive PennyLane Dataset: Curated 3,347 PennyLane-specific code samples from GitHub repositories, quantum computing textbooks, and official PennyLane documentation, each enriched with contextual descriptions and annotations to support instruction tuning.
@@
    Systematic Data Refinement and Annotation: Developed a robust pipeline for dataset cleaning, metadata enrichment, and conversion of code snippets into instruction–query pairs using the GPT-4o API, ensuring compatibility with LLM training workflows.
@@
    RAG-Based Evaluation Framework: Designed a RAG evaluation framework to assess the dataset’s impact on model performance. Evaluated GPT-4o mini, Claude 3.5, Qwen 7B, and LLaMa 4 Maverick 17B under both retrieval-augmented and non-augmented settings, scoring outputs on functional correctness.
Summary of Key Results: Retrieval augmentation with PennyLang yields substantial relative gains for smaller open-source models: Qwen 7B’s success rate rises from 8.71% to 41.66% ($\approx$32.95% relative increase), and LLaMa 4 improves from 78.78% to 84.84% ($\approx$7.7% relative increase). In contrast, stronger baselines (GPT-4o-mini and Claude 3.5 Sonnet) do not benefit from naive full-context retrieval, their best performance is achieved without retrieval (85.60% and 95.07%, respectively), with full retrieval resulting in slight degradation. Detailed analysis is presented in Section 5.
@@ II Background and Related Work
TABLE I: Summary of Qiskit and PennyLane Datasets for LLM Training
Dataset Name Framework Source Samples Available Dataset Type Format Qiskit Framework Datasets Qiskit Textbook Dataset [32] Qiskit Qiskit.org Textbook Multiple Structured Qiskit Code Python Scripts IBM Quantum Challenge Data [33] Qiskit IBM Quantum Experience Derived Instruction-Response Format JSON/Notebook Qiskit Code Repository [34] Qiskit GitHub (Qiskit AI repos) Thousands Raw Qiskit Code Samples Python Scripts Q-Gen Quantum Circuit [35] Qiskit Kaggle (Possible) 309 Pretrained Benchmark Data CSV/JSON QDataSet [36] Qiskit GitHub 52 Quantum ML Data Various QCircuitNet [37] Qiskit Research Paper 4,800 Quantum Algorithm Design Benchmark Python/QASM PennyLane Framework Datasets PennyLane Documentation [38] PennyLane PennyLane.ai Docs Various Raw PennyLane Code Samples Python Scripts PennyLane GitHub Repos [39] PennyLane GitHub Open-Source Scattered Unstructured Quantum Code Python Scripts PennyLang (Ours) PennyLane Open-Source Dataset 3,347 Curated and Structured Dataset JSON file
The intersection of LLMs and quantum computing is an emerging research domain. While LLMs have demonstrated impressive capabilities in NLP [40] and automated code generation [41], their application in quantum software development remains underexplored. Most existing studies focus on classical programming languages or quantum frameworks such as Qiskit [22], with little attention to the PennyLane framework [19], despite its popularity in hybrid quantum-classical workflows. This section presents an in-depth review of related work in three key areas:
@@ II-A LLMs for Code Generation in Classical Programming
Transformer-based language models have revolutionized automated code generation, significantly advancing classical software development. OpenAI’s Codex [26], trained on large-scale public code repositories, enabled natural language-to-code translation, setting a new standard for programming assistants. CodeBERT [42] further improved source code understanding through bidirectional transformers, while InCoder [29] introduced multi-turn interactive capabilities for code infilling and synthesis. AlphaCode [25] demonstrated that LLMs could generate solutions for competitive programming problems, reaching near-human performance. To rigorously evaluate these models, benchmarks such as HumanEval [24], MBPP [43], and CodeXGLUE [44] have been proposed, each offering complementary perspectives on correctness, diversity, and execution fidelity. These benchmarks collectively emphasize the importance of well-curated, structured datasets in improving LLM-driven code generation.
TABLE II: Comparison: Related Work & Our Contributions
Research Area Existing Works Our Contributions LLMs for Code Generation Codex [26] CodeBERT [42] PennyLane-specific dataset LLMs for Quantum Programming Qiskit Assistant [22] [45] Hybrid quantum- classical dataset Quantum Code Datasets QCircuitNet [37] QDataset [37] First PennyLane dataset
@@ II-B LLMs for Quantum Programming
In comparison, quantum programming has seen limited adoption of LLMs, with most efforts focused on the Qiskit framework. IBM’s Qiskit Code Assistant [22] supports Qiskit programs’ debugging and optimizations, while [45] adapted BERT-based models for enhanced quantum code comprehension. CodeLlama-Quantum [27] extended Meta’s CodeLlama to address Qiskit-specific use cases. Despite these advances, key limitations persist. The field lacks large, high-quality datasets for training and evaluation, and quantum circuits demand deeper reasoning due to entanglement, measurement, and complex control logic. Furthermore, hybrid programming paradigms like PennyLane interleave classical and quantum instructions, requiring models to understand multi-modal code patterns and contextual dependencies. To address these challenges, we propose PennyLang, the first PennyLane-centric dataset, supported by an evaluation pipeline that integrates RAG to improve code relevance and contextual grounding. Our contributions advance AI-assisted quantum programming by enabling LLMs to provide high-quality, domain-specific assistance in PennyLane.
Figure 2: Methodology for PennyLang Quantum Code Generation and Evaluation. We collect and refine quantum code from GitHub, textbooks, and PennyLane documentation, then use GPT-4o to convert cleaned snippets into instruction–query pairs (with code and tests). These examples are embedded in Chroma, and LangChain performs MMR-based retrieval to produce RAG-augmented versus vanilla prompts. GPT-4o-mini, Claude 3.5 Sonnet, LLaMa 4-17B Maverick, and Qwen 2.5-7B-Instruct-Turbo are evaluated on functional correctness to quantify the benefit of PennyLang dataset.
@@ III-A1 Data Collection from GitHub Repositories
To systematically gather PennyLane-specific code, we utilized the GitHub API to search for publicly accessible repositories with permissive open-source licenses (e.g., MIT, Apache 2.0, BSD). The selection criteria required that the term "PennyLane" be present in either the repository name, description, or README file. To maintain dataset quality and avoid redundancy, we followed these constraints:
@@
  • Only Python (.py) files were extracted for further analysis.
@@
Since PennyLane maintains backward compatibility and rarely deprecates features [19], we did not impose a time constraint on data collection, ensuring a broad dataset scope. Additionally, introductory comment blocks containing license information, author details, or unrelated metadata were removed to retain only executable and relevant code. This curation process resulted in a high-fidelity dataset of PennyLane-specific implementations. In total, 1,321 samples were collected from the official PennyLaneAI GitHub repository [39].
@@ III-A2 Extracting Quantum Code from Books
In addition to GitHub, we extracted PennyLane-related code snippets from two quantum computing books, Quantum Machine Learning: An Applied Approach  [46] and A Practical Guide to Quantum Machine Learning and Quantum Optimization [47]. These books provide theoretical foundations and applied examples of PennyLane-based quantum machine learning (QML) and variational quantum algorithms (VQAs). The extraction process involved:
@@
    Manually verification with explanations from the books, converting descriptive text into structured comments.
@@ III-A3 Scraping Code from PennyLane Documentation
The official PennyLane documentation [38] is a comprehensive resource covering topics such as quantum circuits, variational algorithms, operators and templates, hybrid quantum–classical optimization, and quantum chemistry applications. To enrich our dataset, we systematically extracted all code snippets from the official website, including those embedded in tutorials and API references. Documentation text accompanying each example was manually converted into inline comments, ensuring every sample is self-contained and context-rich. This approach improves LLM interpretability, enabling models trained on the dataset to produce both functional code and explanatory text.